Three resistorsof values 2 Ω, 6 Ω and 12 Ω are connected across a DC voltage source as shown in Fig. 26-17. If the total current through the circuit is 2.0 A, what is the appliedvoltage?

a) 3.0 V
b) 1.5 V
c) 2.0 V
d) 6.0 V
e) 2.7 V
Answer #1

1/R=1/r1+1/r2+1/r3

=1/2+1/6+1/12

=9/12

R=12/9Ω

i=2 A

V=i*R

=2*9/12

=1.5 V

ans B

answered by: Nenny
Answer #2
total current
i=v/(2||6||12)
i=v/(4/3)
v=4/3*2
v=8/3 volts
answered by: abdul